A 9.00 kg mass is moving to the right with a velocity of 14.0 m/s. A 12.0kg mass is moving to the left with a velocity of 5.00 m/s. Assuming that these two balls have a head on collision and stick together, what will be the final velocity of the combination?
1
Expert's answer
2018-12-13T11:05:37-0500

Answer on Question 83764, Physics, Other

Question:

A 9.0 kg mass is moving to the right with a velocity of 14.0 m/s. A 12.0 kg mass is moving to the left with a velocity of 5.0 m/s. Assuming that these two balls have a head on collision and stick together, what will be the final velocity of the combination?

Solution:

We can find the final velocity of the combination from the law of conservation of momentum:

"m_1v_1+m_2v_2=(m_1+m_2)v_f,"

here,

"m_1=9.0kg"

is the mass moving to the right,

"m_2=12.0 kg"

is the mass moving to the left,

"v_1=14.0 m\/s"

is the velocity of the mass moving to the right,

"v_2=5.0 m\/s"

is the velocity of the mass moving to left,

"v_f"

is the final velocity of the combination.

Let's choose the direction to the right as the positive. Then, we can write:

"m_1v_1-m_2v_2=(m_1+m_2)v_f,""v_f=(m_1v_1-m_2v_2)\/(m_1+m_2)."

Finally, let's substitute the numbers:

"v_f=(9.0kg * 14.0m\/s - 12.0kg * 5.0m\/s)\/(9.0kg + 12.0 kg)=3.14m\/s."

Answer:

"v_f=3.14m\/s"

.